OpenStudy (egbeach):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form

OpenStudy (xapproachesinfinity):

the unit vector is found this way \(\Huge \vec{a}=\frac{\vec{u}}{||\vec{u}||}\) where is a is our unit vector

OpenStudy (xapproachesinfinity):

||u|| is the magnitude of u found by this \(\Huge ||\vec{u}||=\sqrt{(-4)^2+(-3)^2}\)

OpenStudy (xapproachesinfinity):

So \(\huge \vec{a}=\frac{<-4,-3>}{\sqrt{(-4)^2+(-3)^2}}\)

OpenStudy (xapproachesinfinity):

we just need to simplify that and done!

OpenStudy (egbeach):

Thank you so much! @xapproachesinfinity

OpenStudy (xapproachesinfinity):

np!

OpenStudy (xapproachesinfinity):

i guess you can do the rest yes?

OpenStudy (egbeach):

yep!